OpenStudy (anonymous):

if (3,-y) is equidistant rom (3,6) and (7,2) then what is y?

OpenStudy (kittiwitti1):

Use the Distance Formula.\[d = \sqrt{\left( x _{2}-x _{1} \right)^2 - \left({y _{2} - y_{1}} \right)^2}\]

OpenStudy (anonymous):

teach me step by step please i dont get it :(

OpenStudy (kittiwitti1):

You find the distance between (3,6) and (7,2), then divide that by 2 for d and use this d to find the distance between either (3,6) and (3,-y) or (3,-y) and (7,2). Or, you could use the Midpoint Formula, if you want.\[m = \left(\frac{ x_1 + x_2 }{ 2 }, \frac{ y_1 + y_2 }{ 2} \right)\]

OpenStudy (nirmalnema):

use distance formula... rt.or (x1-x2)^2+(y1-y2)^2 (3-3)^2+(6+y)^2=(3-7)^2+(2+y)2 sq rt can be removed by squaring both the side

OpenStudy (anonymous):

ahh.. hard

OpenStudy (kittiwitti1):

Actually the distance formula doesn't work, and you need to try before you say it's too hard @jacalneaila \[d = \sqrt{(2-6)^2 + (7-3)^2} = \sqrt{16 + 16} \approx 5.66\]Midpoint Formula it is, then. \[m = \left( \frac{ x_1 + x_2 }{ 2 }, \frac{ y_1 + y_2 }{ 2 } \right) = \left( \frac{ 3+7 }{ 2 }, \frac{ 6+2 }{ 2 } \right)\]Can you figure out the rest?

OpenStudy (anonymous):

that solutio??n

OpenStudy (kittiwitti1):

It's part of it.

OpenStudy (anonymous):

where the y??

OpenStudy (adilalvi):

y=8

OpenStudy (kittiwitti1):

.-. m = (x,y)

OpenStudy (kittiwitti1):

@adilalvi give her a chance to work it out herself

OpenStudy (adilalvi):

ok

OpenStudy (anonymous):

how

OpenStudy (kittiwitti1):

\[m = \left( \frac{ 10 }{ 2 }, \frac{ 8 }{ 2 } \right) = \left( 5, 4 \right)\]Find the distance between (5,4) and one of the original given points, then use that to find the y value

OpenStudy (adilalvi):

-y+6+2=0 -y+8=0 y=8

OpenStudy (kittiwitti1):

I've already listed out the Distance Formula for you, so work it out yourself and I'll tell you whether you got it right or not.

OpenStudy (nirmalnema):

further simplifing through distance formula 36+12y+y^2=16+4y++4+y^2 y^2 can be cancelled out from both the sides. 36+12y=20+4y 16=-8y y=-2

OpenStudy (anonymous):

huh..going crazy right now i dnt understand

OpenStudy (kittiwitti1):

-sigh- Hold on, I'll type it out for you.

OpenStudy (anonymous):

what? O.o

OpenStudy (kittiwitti1):

I just realized I did the Distance Formula wrong in the first part lol\[\sqrt{(2-4)^2 - (7-5)^2} = \sqrt{4-4} = 0 \]Hmm.. that's strange.

OpenStudy (kittiwitti1):

You know it would be better if you graphed it out before solving.

OpenStudy (anonymous):

what O.o

OpenStudy (kittiwitti1):

Graph it out.

OpenStudy (kittiwitti1):

Honestly I think @nirmalnema is right, but I just wanna check .-.

OpenStudy (anonymous):

help please :(

OpenStudy (nirmalnema):

@jacalneaila you ask what is not clear to you.?

OpenStudy (anonymous):

all :(

OpenStudy (anonymous):

teach me step by step

OpenStudy (nirmalnema):

(3,-y) is equidistant rom (3,6) and (7,2) then what is y? x1=3 , 7 x2=3 y1=2 , 6 y2= it is unknown.. you have to find it...

OpenStudy (anonymous):

how?

OpenStudy (kittiwitti1):

Ask @AllTehMaffs :p

OpenStudy (nirmalnema):

just consider it so that you can easily put the values in distance formula

OpenStudy (anonymous):

how teach me please for my exam tommorow

OpenStudy (anonymous):

(for reference)  So do you know what equidistant means?

OpenStudy (anonymous):

(the unknown point is not correct in that drawing)

OpenStudy (anonymous):

@jacalneaila ? Do you know what "equidistant" means?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

"equidistant" means "equal in distance" So saying (3,-y) is "equidistant" from (3.6) and (7,2) means that the distance between (3,-y) and (3,6) is the same as the distance between (3,-y) and (7,2). Does that make sense?

OpenStudy (anonymous):

So the line segment that connects points (3,-y) and (3,6) is the same length as the line segment that connects points (3,-y) and (7,2).

OpenStudy (anonymous):

Do you know what (3,6) means? Can you plot (3,6) on a graph?

OpenStudy (anonymous):

huh.. im going crazy right now

OpenStudy (anonymous):